# Hilbert polynomials of Calabi Yau hypersurfaces in toric varieties and lattice points in polytope boundaries

### v1

@@ 1.2. Calabi-Yau hypersurfaces in toric varieties
As a Corollary of Theorem 1, the combinatorial computations in Section 3.3 of [22] of lattice point counts in boundaries of low dimensional simplices give the following computations of Hilbert polynomials of low dimensional Calabi Yau manifolds: [^3] In these examples the Hilbert polynomials were computed by counting lattice points by inspection for small $k,$ and deducing the coefficients of the Hilbert polynomial from that data. See [22].
@@ 3. The Euler class in $K$-theory
$\bigotimes_{i=1}^{n}L_{i}^{*}=\bigotimes_{i=1}^{n}(1-(1-L_{i}^{*}))$
$\sum_{\ell=0}^{n}(-1)^{\ell}\sum_{\begin{subarray}{c}I\subset\{1,\ldots,d\}\\ |I|=\ell\end{subarray}}\bigotimes_{i\in I}(1-L_{i}^{*})$
@@
$\prod_{i=1}^{n}x_{i}=\prod_{i=1}^{n}(1-(1-x_{i}))=\sum_{\ell=0}^{n}(-1)^{\ell}\sum_{\begin{subarray}{c}I\subset\{1,\ldots,d\}\\ |I|=\ell\end{subarray}}\prod_{i\in I}(1-x_{i})$
@@
Let $I\subset\{1,\ldots,d\}$ and let $F_{I}=\bigcap_{i\in I}F_{i}$ be a face of $\Delta$. If $\Delta$ is Delzant, so is $F_{I}$. In particular, the divisor $D_{I}=\cap_{i\in I}D_{i}$ is smooth.
This lemma will follow from the following lemma by induction on codimension.
@@
Suppose $\Delta$ is a Delzant polytope and let $F$ be a facet of $\Delta$. Then $F$ is Delzant.
@@
Let $F\subset\Delta$ be a facet and suppose $F$ is not Delzant.
Then there exists a vertex $v$ of $F$ (hence also of $\Delta$) and primitive edge vectors $n_{1},\ldots,n_{m-1}$ lying in the plane $P\subset\mathbb{R}^{m}$ containing $F$ with
(4.3) $|n_{1}\wedge\cdots\wedge n_{m-1}|>1.$
Let $n_{m}$ be the unique primitive edge vector at $v$ to $\Delta$ at $v$ not lying in $P$. Then
$|n_{1}\wedge\cdots\wedge n_{m-1}\wedge n_{m}|>1$
which is not possible, since $\Delta$ is Delzant.∎
Remark. Alternatively, $D_{I}=\mu^{-1}(F_{I})$ is a component of the fixed point set of a subtorus $S\subset T^{m}$, therefore smooth.
@@ 5. Formulas for lattice points in polytope boundaries and the Proof of Theorem 2
$=\int_{M}\operatorname{Td}(T{M})\,e^{c_{1}(L)}\left(1-e^{-\sum_{i=1}^{d}c_{1}(LF_{i})}\right)$
@@
$\operatorname{Td}(T{M})=\prod_{i=1}^{d}\frac{c_{i}(LF_{i})}{1-e^{-c_{i}(LF_{i})}}$
@@
$=\int_{M}\left(\prod_{i=1}^{d}\frac{c_{1}(L_{F_{i}})}{1-e^{-c_{i}(LF_{i})}}\right)e^{c_{i}(L)}\left(1-e^{-\sum_{i=1}^{d}c_{i}(LF_{i})}\right)$
$=\int_{M}\left(\prod_{i=1}^{d}\frac{c_{i}(L_{F_{i}})}{e^{c_{1}(L_{F_{i}})/2}-e^{-c_{1}(L_{F_{i}})/2}}\right)\frac{1}{\prod_{i=1}^{d}e^{-c_{1}(L_{F_{i}})/2}}\left(1-e^{-\sum_{i=1}^{d}c_{1}(LF_{i})}\right)e^{c_{1}(L)}$

### v2

@@ 1.2. Calabi-Yau hypersurfaces in toric varieties
As a Corollary of Theorem 1, the combinatorial computations in Section 3.3 of [22] of lattice point counts in boundaries of low dimensional simplices give the following computations of Hilbert polynomials of low dimensional Calabi Yau manifolds. This can be an efficient way of computing the Hilbert polynomials: In these examples, the Hilbert polynomials were computed by counting lattice points by inspection for small $k,$ and deducing the coefficients of the Hilbert polynomial from that data. See [22] for details.
@@ 3. The Euler class in $K$-theory
$\bigotimes_{i=1}^{n}L_{i}^{*}=\bigotimes_{i=1}^{n}(1-(1-L_{i}^{*}))=\sum_{\ell=0}^{n}(-1)^{\ell}\sum_{\begin{subarray}{c}I\subset\{1,\ldots,d\}\\ |I|=\ell\end{subarray}}\bigotimes_{i\in I}(1-L_{i}^{*}),$
@@
$\prod_{i=1}^{n}x_{i}=\prod_{i=1}^{n}(1-(1-x_{i}))=\sum_{\ell=0}^{n}(-1)^{\ell}\sum_{\begin{subarray}{c}I\subset\{1,\ldots,d\}\\ |I|=\ell\end{subarray}}\prod_{i\in I}(1-x_{i}).$
@@
Let $I\subset\{1,\ldots,d\}$ and let $F_{I}=\bigcap_{i\in I}F_{i}$ be a face of $\Delta$. The divisor $D_{I}=\cap_{i\in I}D_{i}$ is a smooth complex submanifold of $M.$
This lemma will follow from the following lemma by induction on codimension, since the divisor $D_{i}$ corresponding to a facet of $\Delta$ is itself a toric variety.
@@
Let $F_{i}$ be a facet of $\Delta$. Then $D_{i}$ is a smooth complex submanifold of $M.$
@@
For each facet $F_{i}$ of $\Delta,$ there exists a codimension-one subtorus $S\subset T^{m}$ of $T^{m},$ so that $D_{i}$ is a component of the fixed set of $S.$ Since the action of $T^{m}$ is holomorphic, so is the action of $S.$ So $D_{i}$ is a smooth complex submanifold of $M.$
@@ 5. Formulas for lattice points in polytope boundaries and the Proof of Theorem 2
$=\int_{M}\operatorname{Td}(T{M})\,e^{c_{1}(L)}\left(1-e^{-\sum_{i=1}^{d}c_{1}(L_{F_{i}})}\right)$
@@
$\operatorname{Td}(T{M})=\prod_{i=1}^{d}\frac{c_{i}(L_{F_{i}})}{1-e^{-c_{i}(L_{F_{i}})}}$
@@
$=\int_{M}\left(\prod_{i=1}^{d}\frac{c_{1}(L_{F_{i}})}{1-e^{-c_{i}(L_{F_{i}})}}\right)e^{c_{i}(L)}\left(1-e^{-\sum_{i=1}^{d}c_{i}(L_{F_{i}})}\right)$
$=\int_{M}\left(\prod_{i=1}^{d}\frac{c_{i}(L_{F_{i}})}{e^{c_{1}(L_{F_{i}})/2}-e^{-c_{1}(L_{F_{i}})/2}}\right)\frac{1}{\prod_{i=1}^{d}e^{-c_{1}(L_{F_{i}})/2}}\left(1-e^{-\sum_{i=1}^{d}c_{1}(L_{F_{i}})}\right)e^{c_{1}(L)}$